How do I solve a problem saying "By law, the maximum slope of a wheelchair ramp is 1/12. " A ramp is designed that is 4 feet high and has a horizontal length of 50 feet. Does this ramp meet the law?" Please explain this to me I am truly confused!!!
Thank you for your time and help!

Answer by Alan3354(69443)

How do I solve a problem saying "By law, the maximum slope of a wheelchair ramp is 1/12. " A ramp is designed that is 4 feet high and has a horizontal length of 50 feet. Does this ramp meet the law?" Please explain this to me
-------------
It's, "Is 4/50 <= 1/12" ??
4/50 = 24/300
1/12 = 25/300
So it's less, and complies with the law.